(1 point) Find the eigenvalues and eigenvectors of the matrix A = | -1 (-13 5 -3 11 = , vi = and t2 = ,02 =
Answer #1

A - 57 L-13 -3 J Eigen valuer TA-ND1-0 1 = 0 - 3 -3- (-1-d) (-3-d) + (13) (9) -0 = d+48 +6850 d = -2781 da = -2-si Eigenrectofor de -21tis digo vectos ( for d = 2 + 1 eigen vector -1-87 -1- for a=-2-si - - 5 | 1 +81 L-13 -3- L -Itri 63* (**)A (: 2,2